OpenStudy (anonymous):

Find the equation of the line passing through (6, -4) and parallel to the line with the equation 2x - 3y - 1 = 0.

OpenStudy (anonymous):

Parallel means same slope

OpenStudy (anonymous):

2x - 3y - 1 2x-3y=1 -3y=-2x+1 y=2/3x-1/3 So the slope is 2/3

OpenStudy (anonymous):

Let's use point-slope form your point is (6,-4) y-y1=m(x-x1) So let plug it in y-(-4)=2/3(x-6)

OpenStudy (anonymous):

Being helpful and answering my question, unlike you who is just filling my question with useless sapm Go ahead, imranmeah91

OpenStudy (anonymous):

*spam

OpenStudy (anonymous):

y-(-4)=2/3(x-6) y+4=2/3(x-6) y+4=2/3x -4 y=2/3x -8

OpenStudy (anonymous):

I have 4 options: A: 2x - 3y - 24 = 0 B: 3x + 2y - 10 = 0 C:2x - y - 16 = 0 D: 2x - 3y - 18 = 0

OpenStudy (anonymous):

y=2/3x -8 Multiply the whole thing by 3 3y=2x-24 2x-3y-24

OpenStudy (anonymous):

Thankyou very VERY much!